# Sequential estimation of disturbed aerodynamic flows from sparse measurements via a reduced latent space

### v1

@@ 2.4 Learning surrogate models in the latent space
As described in Section 2.3, the two essential components for the sequential filter are the forecast and observation operators. We have already learned a surrogate observation operator, $\boldsymbol{p}=\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})$, by virtue of the side network in our physics-augmented autoencoder architecture. With the low-dimensional latent states $\boldsymbol{\xi}$ established via the autoencoder, we now aim to learn the surrogate model that governs the temporal evolution of the flow state in this reduced latent space. These surrogate models replace traditional high-fidelity solvers by approximating the forecast and observation operators directly in the reduced space, enabling fast and efficient DA. For a comprehensive overview of various approaches to learning surrogate models for dynamical systems, readers are referred to Bach et al., (2024).
Figure 4: Bifurcation in learned latent trajectories for various gust-encounter aerodynamic cases, with each trajectory shown in a different color. For each case, the latent trajectory deviates from its corresponding undisturbed path (base flow with the airfoil at $\alpha=20^{\circ}$) at the gust onset time $t_{o}$, reflecting the system’s response to external disturbance.
@@
where $\boldsymbol{f}_{\boldsymbol{W}}\left(\boldsymbol{\xi}_{k-1}\right)=\tilde{\boldsymbol{f}}_{\boldsymbol{W}}\left(\boldsymbol{\xi}_{k-1}\right)\Delta t+\boldsymbol{\xi}_{k-1}$ with $k$ the time index.
A schematic illustration of latent trajectories under different random gust conditions is shown in Figure 4. Prior to the introduction of the gust, all trajectories closely follow the undisturbed path, indicating that the system initially behaves as in the baseline case. For simplicity in the illustration in Figure 4, the baseline is flat, corresponding to a steady flow with a time-invariant point in the latent space. However, most cases considered in this work will have time-periodic baseline trajectories reflecting the bluff-body vortex shedding of a wing at a significant angle of attack. When the gust is introduced at a randomly selected time (denoted by solid circles in the figure) within the shedding cycle, the corresponding trajectory deviates from the baseline and subsequently returns to the baseline path once the gust has convected sufficiently far from the airfoil. This bifurcation behavior captures the transient influence of the gust on the flow dynamics. The purpose of the forecast model is not to predict the point of bifurcation from the baseline (a task that is assigned to the filter’s measurement update), but rather, to capture the subsequent dynamics of the flow in the latent space. To ensure the Neural ODE accurately models these dynamics and to provide diverse initial conditions across different aerodynamic cases, snapshots for each disturbed case are restricted to begin 10 time steps ($\Delta t$) after the gust onset and continue for 200 snapshots, encompassing the full gust–airfoil interaction. This targeted training strategy enables the Neural ODE to learn the disturbed dynamics effectively, ensuring that during sequential filtering, it can faithfully track the perturbed latent trajectories shortly after the measurement-corrected state aligns with the true disturbed path. The network architecture is listed in Table 2 and is trained on both undisturbed and disturbed cases.
@@
where $\beta_{f,1}$ and $\beta_{f,2}$ are weighting factors assigned to the rollout and one-step losses, respectively. To emphasize accurate prediction during the initial phase of gust–airfoil interaction, we decay the rollout weighting over time as $\beta_{f,1}=1-0.5(t-1)/(T-1)$ for $t=1,\cdots,T$, with $T$ denoting the final rollout time step. This schedule ensures that higher weight is given to early predictions during disturbance events and gradually reduces the contribution of later predictions within the rollout horizon. In this formulation, we set $\beta_{f,2}=150$, as it was observed that these weights give better predictions.
With the surrogate models now learned in the reduced-order space, all necessary components for sequential DA are in place.
@@ 3 Results and discussion
In this section, we first focus on the compression of full-dimensional flow data into a reduced-dimensional latent space via a physics-augmented autoencoder. Then, we discuss the training of surrogate models necessary for filtering. Finally, we present results for data assimilation in the latent space.
@@ 3.1 Construction of latent space and learned operators
Previous researchers have demonstrated that high-dimensional disturbed aerodynamic flows can be effectively projected onto very low-dimensional manifolds, as low as three-dimensional, using nonlinear autoencoders (Fukami and Taira, 2023, Mousavi and Eldredge, 2025, Fukami et al., 2025). However, it is important to note that when additional regularization terms—such as temporal smoothness and pressure reconstruction losses—are incorporated into the training objective, the dimensionality of the latent space must generally be increased to compensate for the stricter constraints imposed on the encoded representation. This increase provides the necessary degrees of freedom to adequately capture the dominant flow features and enhance their observability within the compressed subspace.
@@
To determine an appropriate latent-space dimension $n$, we take the following systematic approach, guided both by the error trends in training the autoencoder and the effective state rank of the sequential filtering in the LREnKF. We first conducted a systematic analysis using the network architecture shown in figure 3 and table 1. Specifically, we examined the validation loss as a function of latent dimension and observed that the loss plateaued beyond $n=7$, indicating diminishing returns in reconstruction fidelity with further increases in dimensionality. This finding is corroborated by a complementary experiment in which the dataset was compressed into a 10-dimensional latent space, followed by sequential DA. As illustrated in figure 5, rarely did more than seven directions in the latent space exhibit sensitivity to variations in the sensor measurements, further supporting the choice of $n=7$. While it is possible to adopt a higher-dimensional latent representation, the goal in this work is to encode the high-dimensional flow field into the most compact latent space that still preserves the information necessary for accurate state estimation. Conversely, selecting a latent dimension smaller than seven risks excessive compression, compromising the uniqueness and identifiability of the latent representation of flow trajectories required for robust sequential filtering and inference. Overall, the choice of latent dimension reflects a tradeoff among reconstruction accuracy, compactness of the representation, and degree of identifiability required for reliable sequential estimation. Based on these analyses, we set $n=7$ as the latent dimension in this study.
@@
Figure 5: The history of state rank $r_{\boldsymbol{\xi}}$ for $n=10$. The flow condition is ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$).
@@
To explore the extent to which the latent dynamics are interpretable, figure 6 presents the histories of the latent vector components for several independent disturbed cases at $\alpha=40^{\circ}$. The undisturbed flow over the airfoil at a high AoA (such as the case in this figure represented by black curves) exhibits characteristic periodic behavior, which is also clearly reflected in the latent space representation. The time window shown spans slightly more than two full vortex shedding cycles. The bifurcation highlighted earlier in figure 4 is evident here as well, manifesting as divergence in the latent component histories due to the gust encounters at random phases of the shedding cycle. While these trajectories implicitly encode the influence of gust characteristics, vortex dynamics, and their interactions, the mapping between individual latent variables and specific physical phenomena remains difficult to disentangle. Determining which components of the latent vector correspond to particular aspects of the flow or external disturbances would require targeted analyses beyond the scope of this work. A systematic study aimed at interpreting these latent representations and rigorously connecting them to the underlying flow physics (e.g. Smith et al., (2024)) remains an important direction for future research.
@@
Dipole gusts consist of a positive and a negative vorticity lobe convecting across the airfoil. Their aerodynamic impact can be understood in terms of their interaction with the primary vortical structures of the separated shear layer. In all investigated cases, a consistent aerodynamic pattern emerges in response to dipole gusts. As the positive lobe of the dipole interacts with the primary negative LE vortex, the effective angle of attack increases transiently, elevates circulation, which produces a sharp lift surge. Shortly thereafter, the negative lobe arrives and disrupts the vortex structure, leading to a rapid drop in lift.
@@
Now that we have demonstrated successful state compression in tandem with an accurately learned observation operator, there remains one more verification step before proceeding to DA: assessing the performance of the learned forecast operator described in Section 2.4. Figure 8 illustrates the history of one predicted latent variable, $\xi_{2}$, during gust passage in one of the test cases. Starting from the true initial state, the autoregressive prediction initially tracks the true trajectory but eventually diverges. At first glance, this deviation may seem unexpected, given that the model was explicitly trained with a rollout loss designed to prevent error accumulation over long prediction horizons. It is important to emphasize, however, that the one-step prediction remains highly accurate, as is essential for DA applications. Indeed, the learned forecast operator provides reliable short-term estimates during rollout predictions, ensuring that sequential updates in filtering can be based on precise time-local dynamics, even if longer-term rollout trajectories diverge. Despite the larger prediction error observed in this example, the inclusion of the rollout loss does stabilize the network: the error does not grow without bound over time. Rather, the discrepancy arises because, while the original high-dimensional dynamics are Markovian, this quality is not necessarily preserved when compressed into a low-dimensional latent space. In other words, two nearby initial conditions in the compressed representation can evolve into significantly different trajectories. This important limitation has been rigorously established in the foundational work of Zwanzig, (1961) and Mori, (1965). According to Mori–Zwanzig theory, the evolution of any reduced subspace can be decomposed into a Markovian term, a memory (history-dependent) term, and an unobservable (noise) term. As demonstrated by Ruiz et al., (2024), explicitly modeling the memory term can significantly improve predictive performance. Following this insight, one promising way to improve the forecast operator is to incorporate history dependence using memory-based NNs such as RNNs. However, this approach comes at the cost of breaking the Markovian assumption required by standard sequential filtering methods. Indeed, the philosophy of our current approach is that the responsibility for correcting deviations lies with the measurement update of each filter step. Thus, we deliberately choose to retain this imperfect, but short-term-accurate, forecast operator to evaluate the performance of our filtering framework under the realistic scenario of a non-ideal predictive model.
@@
Figure 8: Example of the history of true (black) and predicted (blue, from the learned forecast operator) latent state for the airfoil at $\alpha=20^{\circ}$ during a gust encounter.
@@ 3.2 Sequential filtering examples
Now, everything is in place to set up the DA task. Pressure measurements from 11 sensors uniformly mounted on the airfoil (see figure 2) in CFD simulations with added noise are used for state estimation via the LREnKF framework. Each member of the ensemble of estimated latent states is then propagated through the pre-trained decoder of the network architecture shown in figure 3 to reconstruct both the lift coefficient and the vorticity field. For all cases in this section, a convergence study was performed, and we selected an ensemble size of $N_{e}=200$. The ensemble members were initialized as Gaussian perturbations from the true state, always a latent representation of the undisturbed initial flow about the airfoil, with a variance of $10^{-4}$. Each assimilation step is observed to take as fast as a few milliseconds.
The results are presented in figures 9 and 10 for the smallest and largest AoA considered in this study, namely $\alpha=20^{\circ}$ and $60^{\circ}$, respectively. In these test cases, random disturbances are introduced at random instants within the first cycle. For both cases, the predicted lift coefficient shows excellent agreement with the true values throughout the assimilation horizon. The estimation closely tracks the temporal variations induced by gust disturbances, with uncertainty intervals remaining narrow except during the disturbance passage, when larger deviations naturally occur. This confirms that the assimilation effectively leverages the sparse pressure measurements to correct the latent state estimate in real time, even when the forecast operator alone exhibits significant predictive error.
@@
Figure 9: Estimated lift and vorticity field for disturbed aerodynamics at the condition ($\alpha=20^{\circ}$, $D_{y}=-1.9$, $\sigma=0.1$, $y_{o}=0.22$, $t_{o}=1.5$). The predicted ensemble mean along with the $95\%$ confidence interval is shown for the lift coefficient and the vorticity field.
Figure 10: Estimated lift and vorticity field for disturbed aerodynamics at the condition ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$). The predicted ensemble mean along with the $95\%$ confidence interval is shown for the lift coefficient and the vorticity field.
To further elucidate the relationship between reconstruction error and the underlying observability properties of the estimator, we examine the temporal evolution of both the state rank $r_{\boldsymbol{\xi}}$ and observation rank $r_{\boldsymbol{p}}$ during the assimilation process. The state rank reflects the number of latent directions actively informed by observations and corrected by the measurement updates, whereas the observation rank quantifies the number of independent pressure observation modes that effectively inform the latent variables. The evolution of these ranks for a representative test case at the highest angle of attack, corresponding to figure 10, is shown in figure 11. The figure highlights the distinctly low-rank nature of the estimator: among the 11 available pressure sensors, the number of informative observation directions varies between 4 and 7 over time, while 5 to 7 latent directions are actively updated in the analysis step of the filter. This behavior was consistently observed across all angles of attack and random disturbance realizations. Notably, attempting DA with a full-rank EnKF in this context can lead to over- or under-correction of the state estimates due to spurious correlations and sampling noise, underscoring the necessity of rank truncation for robust and reliable flow reconstruction.
Figure 11: Effective state rank $r_{\boldsymbol{\xi}}$ and observation rank $r_{\boldsymbol{p}}$ for disturbed aerodynamics under the condition ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$).
[…]
Figure 12: The first, second, third, and seventh latent state modes at four distinct time instances during the gust interaction for the case ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$).
As observed in figure 12, the latent directions associated with very small eigenvalues correspond to modes that are effectively unobserved by the pressure sensors. For the case $\alpha=20^{\circ}$, we conduct a dedicated investigation of these unobservable latent directions by explicitly constructing perturbations in the approximate nullspace of the Jacobian of the observation operator, $\nabla\boldsymbol{h}_{\boldsymbol{W}}$. The resulting vorticity reconstructions, shown in figure 13, illustrate that even substantial deviations in the latent state along these directions can produce predicted pressure measurements that remain well within the uncertainty bounds of the sensors, yet correspond to markedly different flow fields. This phenomenon provides a direct demonstration of the many-to-one mapping between latent representations and pressure signatures. Notably, the vorticity differences induced by nullspace perturbations manifest predominantly in regions where the sensors have limited sensitivity: the gust core, the separated shear layers along the suction side, and the wake. This explains the residual ambiguity and elevated uncertainty observed in the reconstructed flow fields in figures 9 and 10. The errors $\varepsilon_{d}$ further highlight that unobservable latent modes can contribute significantly to the overall vorticity prediction error. These results underscore that, in the absence of additional measurement modalities or prior regularization, multiple latent states can remain indistinguishable from the perspective of the available sensors, fundamentally limiting the estimator’s ability to uniquely resolve the true flow configuration. This ambiguity becomes more pronounced once the gust has moved away from the airfoil, a regime in which the pressure sensors exhibit reduced sensitivity to residual flow variations. As a result, the likelihood of multiple plausible latent solutions increases, particularly in regions or time intervals where the flow-induced pressure signals are weak. This issue could be mitigated by incorporating a maximum observability constraint during training the autoencoder detailed in section 2.2. We explored this strategy through a modified formulation that enforces a unique latent representation consistent with the pressure observations, leading to significantly improved accuracy in latent state estimation. However, under this constraint, the vorticity decoder exhibited heightened sensitivity to small perturbations in the latent variables, such that even minor estimation errors resulted in non-physical artifacts and substantial degradation in reconstruction quality. This tradeoff highlights the challenge of balancing observability in measurement space and robustness in high-dimensional state reconstruction when designing compressed latent-variable-based estimators for unsteady flow reconstruction.
Figure 13: The unobservable vorticity plots for disturbed flow at condition ($\alpha=20^{\circ}$, $D_{y}=-1.9$, $\sigma=0.1$, $y_{o}=0.22$, $t_{o}=1.5$).
It is also insightful to examine the observation modes corresponding to the sensor configuration and their role in informing state estimation. Figure 14 presents the leading and trailing observation eigenmodes—specifically $u_{1}$, $u_{2}$, $u_{3}$, and $u_{11}$—with the same flow conditions as figure 12. The sensor indices follow the ordering shown in figure 3, where sensor 1 is located at the trailing edge on the upper surface, and the numbering proceeds counterclockwise around the airfoil, ending at the trailing edge on the lower surface. Each observation mode is a weighted combination of sensor outputs that, through the Kalman gain, most strongly influences the filter correction in a corresponding latent direction. For instance, $u_{1}$ reflects the most informative observation direction. The weight assigned to each sensor in these modes reflects its relative contribution to that direction in the observation space. From these plots, we observe that during the critical time interval $t\in[3.6,4.0]$, when the gust approaches and begins interacting with the airfoil on the pressure side, leading-edge sensors 6 through 9 on the lower surface become the most responsive. These sensors exhibit consistently high contributions across the dominant observation modes, indicating their central role in informing the estimator throughout the gust-airfoil-wake interaction ($t\in[3.6,5.0]$). While the relative contributions of other sensors vary over time, the leading-edge and lower-side sensors remain the most informative across modes and time steps.
It is worth emphasizing that this observation contrasts with findings from our earlier study (Mousavi and Eldredge, 2025), in which the upstream disturbance was modeled using a Taylor vortex. A Taylor vortex typically introduces a stronger and more disruptive disturbance compared to a vortex dipole, leading to intense interaction with the primary shear layer and edge vortices on the suction side. As a result, the suction-side sensors were found to be highly sensitive and informative in that setting. In contrast, the dipole-like gust generated by Gaussian forcing in the present work interacts more locally with the pressure side, producing a more focused and asymmetric sensor response. This distinction highlights the importance of gust structure in shaping the spatial distribution of sensor informativeness and ultimately determining the directions along which the latent states can be effectively constrained.

### v2

@@ 2.4 Learning surrogate models in the latent space
As described in Section 2.3, the two essential components for the sequential filtering are the forecast and observation operators. We have already learned a surrogate observation operator, $\boldsymbol{p}=\boldsymbol{h}_{\boldsymbol{W}}(\boldsymbol{\xi})$, by virtue of the side network in our physics-augmented convolutional autoencoder architecture. With the low-dimensional latent states $\boldsymbol{\xi}$ established via the autoencoder, we now aim to learn the temporal evolution of the flow state in this reduced latent space. These surrogate models replace traditional high-fidelity solvers by approximating the forecast and observation operators directly in the reduced space, enabling fast and efficient DA. For a comprehensive overview of various approaches to learning surrogate models for dynamical systems, readers are referred to Bach et al. [2].
Figure 4: Bifurcation in learned latent trajectories for various gust-encounter aerodynamic cases, with each trajectory shown in a different color. For each case, the latent trajectory deviates from its corresponding undisturbed path (base flow with the airfoil at $\alpha=20^{\circ}$) at the gust onset time $t_{o}$, reflecting the system’s response to external disturbance. All trajectories are obtained directly from the encoded data using the trained autoencoder, without any data assimilation or neural ODE.
@@
where $\boldsymbol{f}_{\boldsymbol{W}}\left(\boldsymbol{\xi}_{k-1}\right)=\tilde{\boldsymbol{f}}_{\boldsymbol{W}}\left(\boldsymbol{\xi}_{k-1}\right)\Delta t+\boldsymbol{\xi}_{k-1}$ with $k$ the time index. Although a Neural ODE is adopted here to model the latent dynamics, the proposed framework is not restricted to this choice. Any alternative dynamical model capable of propagating the latent state forward in time can be incorporated within the same sequential filtering structure, provided it can describe this dynamical update at a wide variety of states. An RNN-based approach is also likely suitable, for example.
[…]
Conditioned on the encoded angle of attack (selected from five discrete AoA values) and the initial latent state, the forecast model is constructed to evolve the flow dynamics during the approach and interaction of a disturbance with the fixed-angle airfoil. The network architecture is summarized in Table 2 and is trained using both undisturbed and gust-disturbed cases.
@@
where $\beta_{f,1}$ and $\beta_{f,2}$ are weighting factors assigned to the rollout and one-step losses, respectively. To emphasize accurate prediction during the initial phase of gust–airfoil interaction, we decay the rollout weighting over time as $\beta_{f,1}=1-0.5(t-1)/(T-1)$ for $t=1,\cdots,T$, with $T$ denoting the final rollout time step. This schedule ensures that higher weight is given to early predictions during disturbance events and gradually reduces the contribution of later predictions within the rollout horizon. In this formulation, we set $\beta_{f,1}=1.0$ and $\beta_{f,2}=150.0$, as it was observed that these weights give better predictions.
With the surrogate models, as well as the process noise and observation noise covariances, now learned in the reduced-order space, all necessary components for sequential DA are in place. Note that after evolving ensembles forward through the deterministic learned forecast operator $\boldsymbol{f}_{\boldsymbol{W}}$, an i.i.d. random process noise term $\boldsymbol{w}_{k}$ is added to each ensemble member to represent model uncertainty. The measurement update acts as a mechanism that identifies departures from the undisturbed baseline trajectory and redirects the state estimate toward the correct disturbed branch. Thus, to ensure timely detection of disturbance onset, data assimilation is performed at every time step, yielding 500 assimilation steps per case over approximately two vortex-shedding cycles (about 250 steps per shedding period). The time step $\Delta t=0.02$ is used consistently for both state evolution and measurement assimilation. Notably, this assimilation frequency remains lower than the operating frequency of practical pressure sensors, underscoring the feasibility of the proposed framework.
@@ 3 Results and discussion
In this section, we first focus on the compression of full-dimensional flow data into a reduced-dimensional latent space via a physics-augmented convolutional autoencoder. Then, we discuss the training of surrogate models necessary for filtering. Finally, we present results for data assimilation in the latent space.
@@ 3.1 Construction of latent space and learned operators
Previous researchers have demonstrated that high-dimensional disturbed aerodynamic flows can be effectively projected onto very low-dimensional manifolds—as low as three-dimensional—using nonlinear autoencoders [15, 32, 14]. However, it is important to note that when additional regularization terms—such as temporal smoothness and pressure reconstruction losses—are incorporated into the training objective, the dimensionality of the latent space must generally be increased to compensate for the stricter constraints imposed on the encoded representation. This increase provides the necessary degrees of freedom to adequately capture the dominant flow features and enhance their observability within the compressed subspace.
@@
To determine an appropriate latent-space dimension $n$, we take the following systematic approach, guided both by the error trends in training the autoencoder and the effective state rank of the sequential filtering in the eigenspace. We first conducted a systematic analysis using the network architecture shown in figure 3 and table 1. Specifically, we examined the validation loss as a function of latent dimension and observed that the loss plateaued beyond $n=7$, indicating diminishing returns in reconstruction fidelity with further increases in dimensionality. This finding is corroborated by a complementary experiment in which the dataset was compressed into a 10-dimensional latent space, followed by sequential DA and tracking the first $r_{\boldsymbol{\xi}}$ eigendirections with $99\%$ cumulative spectral energy. As illustrated in figure 5, rarely did more than seven directions in the latent space exhibit sensitivity to variations in the sensor measurements, further supporting the choice of $n=7$. While it is possible to adopt a higher-dimensional latent representation, the goal in this work is to encode the high-dimensional flow field into the most compact latent space that still preserves the information necessary for accurate state estimation. Conversely, selecting a latent dimension smaller than seven risks excessive compression, compromising the uniqueness and identifiability of the latent representation of flow trajectories required for robust sequential filtering and inference. Overall, the choice of latent dimension reflects a tradeoff among reconstruction accuracy, compactness of the representation, and degree of identifiability required for reliable sequential estimation. Based on these analyses, we set $n=7$ as the latent dimension in this study.
@@
Figure 5: The history of state rank $r_{\boldsymbol{\xi}}$ for $n=10$. The first $r_{\boldsymbol{\xi}}$ eigenvalues of $\boldsymbol{C}_{\boldsymbol{\xi}}$ retain $99\%$ of cumulative energy. The flow condition is ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$).
@@
To explore the extent to which the latent dynamics are interpretable, figure 6 presents the histories of the latent vector components for several independent disturbed cases at $\alpha=40^{\circ}$. The undisturbed flow over the airfoil at a high AoA (such as the case in this figure represented by black curves) exhibits characteristic periodic behavior, which is also clearly reflected in the latent space representation. The time window shown spans slightly more than two full vortex shedding cycles. The bifurcation highlighted earlier in figure 4 is evident here as well, manifesting as divergence in the latent component histories due to the gust encounters at random phases of the shedding cycle. While these trajectories implicitly encode the influence of gust characteristics, vortex dynamics, and their interactions, the mapping between individual latent variables and specific physical phenomena remains difficult to disentangle. Determining which components of the latent vector correspond to particular aspects of the flow or external disturbances would require targeted analyses beyond the scope of this work. A systematic study aimed at interpreting these latent representations and rigorously connecting them to the underlying flow physics (e.g. [40]) remains an important direction for future research.
@@
Now that we have demonstrated successful state compression in tandem with an accurately learned observation operator, there remains one more verification step before proceeding to DA: assessing the performance of the learned forecast operator described in Section 2.4. Figure 8 illustrates the history of one predicted latent variable, $\xi_{2}$, during gust passage in one of the test cases. Starting from the true initial state of a disturbed path, the autoregressive prediction initially tracks the true trajectory but eventually diverges. At first glance, this deviation may seem unexpected, given that the model was explicitly trained with a rollout loss designed to prevent error accumulation over long prediction horizons. It is important to emphasize, however, that the one-step prediction remains highly accurate, as is essential for DA applications. Indeed, the learned forecast operator provides reliable short-term estimates during rollout predictions, ensuring that sequential updates in filtering can be based on precise time-local dynamics, even if longer-term rollout trajectories have larger errors. Despite the larger prediction error observed in this example, the inclusion of the rollout loss does stabilize the network: the error does not grow without bound over time. Rather, the discrepancy arises because, while the original high-dimensional dynamics are Markovian, this quality is not necessarily preserved when compressed into a low-dimensional latent space. In other words, two nearby initial points in the compressed representation can evolve into significantly different trajectories. This important limitation has been rigorously established in the foundational work of Zwanzig [45] and Mori [31]. According to Mori–Zwanzig theory, the evolution of any reduced subspace can be decomposed into a Markovian term, a memory (history-dependent) term, and an unobservable (noise) term. As demonstrated by Ruiz et al. [38], explicitly modeling the memory term can significantly improve predictive performance. Following this insight, one promising way to improve the forecast operator is to incorporate history dependence. However, the philosophy of our current approach is that the responsibility for correcting deviations lies with the measurement update of each filter step. Thus, we deliberately choose to retain this imperfect, but short-term-accurate, forecast operator to evaluate the performance of our filtering framework under the realistic scenario of a non-ideal long-term predictive model. We emphasize that one could also adopt a different strategy—using, e.g., an RNN—and likely achieve similar estimation performance, given the forecast operator’s limited role in this framework. Any other strategy would be similarly challenged by the lack of information about the gust arrival, and thus require the use of frequent measurement updates to detect this arrival.
@@
Figure 8: Example of the history of true (black) and predicted (blue, from the learned forecast operator) latent state for the airfoil at $\alpha=20^{\circ}$ during a gust encounter. Nevertheless, the one-step prediction required by DA remains highly accurate.
@@ 3.2 Sequential filtering examples
Now, everything is in place to set up the DA task. Pressure measurements from 11 sensors uniformly mounted on the airfoil (see figure 2) in CFD simulations with added noise are used for state estimation via the EnKF framework. Each member of the ensemble of estimated latent states is then propagated through the pre-trained decoder of the network architecture shown in figure 3 to reconstruct both the lift coefficient and the vorticity field. For all cases in this section, a convergence study was performed, and we selected an ensemble size of $M=200$. We chose to use a large ensemble size to ensure that ensemble statistics were fully and unambiguously converged. The large ensemble is tractable because of the modest dimensions of the latent and measurement spaces. The ensemble members are initialized from a Gaussian distribution whose mean is intentionally offset from the true latent state corresponding to the undisturbed initial flow over the airfoil. Specifically, the initial ensemble mean is biased by 0.5 relative to the truth, with a variance of 0.25, ensuring that the prior distribution is sufficiently separated from the true state and that convergence is nontrivial. The computational efficiency of the proposed framework is demonstrated by performing 500 sequential forecast–analysis cycles over ten convective time units in approximately two seconds on a single GPU. This corresponds to an average runtime of roughly 4 milliseconds per assimilation cycle, demonstrating that the proposed framework is computationally lightweight and suitable for real-time implementation for practical sensor sampling rates.
Figure 9: Estimated lift and vorticity field for disturbed aerodynamics at the condition ($\alpha=20^{\circ}$, $D_{y}=-1.9$, $\sigma=0.1$, $y_{o}=0.22$, $t_{o}=1.5$). The predicted ensemble mean, along with the $95\%$ confidence interval, is shown for the lift coefficient and the vorticity field.
Figure 10: Estimated lift and vorticity field for disturbed aerodynamics at the condition ($\alpha=60^{\circ}$, $D_{y}=-0.71$, $\sigma=0.12$, $y_{o}=-0.19$, $t_{o}=3.3$). The predicted ensemble mean, along with the $95\%$ confidence interval, is shown for the lift coefficient and the vorticity field.
The results are presented in figures 9 and 10 for the smallest and largest AoA considered in this study, namely $\alpha=20^{\circ}$ and $60^{\circ}$, respectively. In these test cases, disturbances with random parameters (strength, size, and orientation) are introduced at arbitrary instants within the first cycle. For both cases, the predicted lift coefficient shows excellent agreement with the true values throughout the assimilation horizon. Even when the initial ensemble distribution deviates substantially from the true initial state, the filter rapidly converges toward the correct trajectory and subsequently tracks the temporal variations induced by gust disturbances with high fidelity. This confirms the estimator’s insensitivity to the initial condition. The associated uncertainty bounds remain tight throughout the evolution, except during the initial DA phase when the ensemble is introduced to be far from the true state with relatively large variance. This demonstrates that the assimilation effectively leverages sparse pressure measurements to correct the latent state estimate—most critically by identifying and capturing the transition to the disturbed trajectory—even when the forecast operator alone exhibits substantial long-term prediction error and is incapable of anticipating the shift from the undisturbed baseline path to the gust-driven response.